OpenStudy (anonymous):

The math team sold 42 hamburger meals and 15 ribs at the pep rally for a total of $243. At the game that night, they sold 36 hamburger meals and 30 ribs meals for a total of $294. What is the cost of 1 rib meal?

OpenStudy (akashdeepdeb):

Let the cost of one hamburger = h Also, let the cost of one rib = r So you can form one equation from the first sentence. 42h + 15r = 243 [Because we are multiplying their number by the cost of one of the entities to get the total cost of the entire meal] You can get the second equation: 36h + 30r = 294 Solve the 2 equations you have and then obtain the price of 1 rib meal, essentially, the value of 'r'. Getting this? :)

OpenStudy (anonymous):

How am I supposed to solve those equations..

OpenStudy (akashdeepdeb):

There are many methods. Try substitution. That is, find the value of h (or r) in terms of the other variable from one equation. Use that value of h (or r) in the second equation to get the numerical value of h (or r). Let's say we had a + b = 4 a - b = 2 Use the first equation to get 'a'. [a + b = 4 implies that, a = 4 - b] Now use this, a value (which is in terms of the other variable) and substitute this in the second equation: (4 - b) - b = 2 4 - 2b = 2 2b = 2 b = 1 a = 4-b = 4-1 = 3 So here a = 3, b = 1. Using this example try finding the value of h and r in your question. Can you do that? :)